OpenStudy (anonymous):

an airplane flew for 5 hours with a tail wind of 25km/h. the return flight against the same wind took them 5.5 hours. find the speed of the airplane in still air!!!!

OpenStudy (anonymous):

let the speed is x so in first flight the resultant velocity=x+25 in return x-25 the distance is (x+25)*5=(x-25)*5.5 .5x=25*5+25*5.5 x=525km/h